# Towards Architecture-Agnostic Untrained Network Priors for Image Reconstruction with Frequency Regularization

### v2

@@ Abstract
Untrained networks inspired by deep image prior have shown promising capabilities in recovering a high-quality image from noisy or partial measurements, without requiring training data. Their success has been widely attributed to the spectral bias acting as an implicit regularization induced by suitable network architectures. However, applications of such network-based priors often entail superfluous architectural decisions, overfitting risks, and slow optimization, all of which hinder their practicality. In this work, we propose efficient, architecture-agnostic methods for a more direct frequency control over the network priors: 1) constraining the bandwidth of the white-noise input, 2) controlling the bandwidth of the interpolation-based upsamplers, and 3) regularizing the Lipschitz constants of the layers. We show that even with just one extra line of code, the overfitting issues in underperforming architectures can be alleviated such that their performance gaps with the high-performing counterparts can be largely closed despite their distinct configurations, mitigating the need for architecture tuning. This then makes it possible to employ a more compact model to achieve similar or superior performance to larger models with greater efficiency. Our regularized network priors compare favorably with current supervised and self-supervised methods on MRI reconstruction and image inpainting tasks, serving as a stronger zero-shot baseline reconstructor. Our code will be made publicly available.
@@ 1 Introduction
Magnetic resonance imaging (MRI) is a mainstream imaging tool for medical diagnosis. Reconstructing MR images from raw measurements refers to the transformation from Fourier spectrum of the object in $k$-space to image space. Since acquiring full $k$-space measurements is time-consuming, under-sampled $k$-space data are often collected to reduce scan times. Accelerated MRI is thus known as an ill-posed inverse problem that conventionally requires handcrafted priors [31, 27] to mitigate the resulting aliasing artifacts in the output images. While supervised learning methods based on convolutional neural networks (CNNs) have demonstrated better reconstruction quality with fewer measurements, their training relies on paired under-sampled and fully-sampled measurements, which are expensive to acquire and could raise issues on robustness and generalization when the acquisition protocol or anatomy changes [23, 24].
@@
Figure 1: Example results from underperforming architectures with 4×4$\times$4 × under-sampling. Turning the left to the right simply by low-pass filtering the white-noise input via a Gaussian blur kernel, which can be implemented as few as one or two lines of code. SSIM (normal-$\uparrow$) values are reported.
@@
Instead of requiring large-scale datasets for capturing prior statistics, the untrained network inspired by deep image prior (DIP) [45] requires only the corrupted or partial measurements and regularizes the reconstruction solely through the inductive biases of its architecture, making it a representative zero-shot self-supervised reconstruction method for a wide range of natural and medical applications [50, 28, 37]. Concretely, DIP parameterizes the unknown desired image via a neural network and optimizes the network parameters such that the output image transformed by the degradation model matches the acquired measurements. Such parameterization is observed to exhibit high resistance to noise and corruption, acting as a form of implicit regularization. Studies have attributed this property to CNN’s inherent spectral bias – the tendency to fit the low-frequency signals before the high-frequency signals (e.g., noise) [41, 5], where the choice of network architecture is often shown to be critically relevant [5, 30, 1].
@@
Despite its great promise, favorable results obtained from the untrained network prior are contingent upon two critical factors: an optimal architecture for a specific task and an early-stopping strategy to prevent overfitting to the noisy, partial measurements. Furthermore, while being agnostic to domain shifts by optimizing on a per-image basis [48], such reconstruction process is inherently slow, hindering its practicality. While these issues are often intertwined — inappropriate and over-parameterized architectures usually exacerbate overfitting and increase runtime, most prior works tackle these challenges separately. For architectural design, existing methods rely on either handcrafting or utilizing neural architecture search techniques [14, 9, 1, 6, 26, 16]. However, the lack of a consensus on the desired architectural prior often leads to a laborious search. Another line of work is dedicated to preventing overfitting through early-stopping [48, 46], subspace optimization [2] or pretraining-then-fine-tuning [3, 35], which also helps shorten the runtime. Similar in spirit to traditional ways of combating overfitting, these methods aim to use fewer trainable parameters with little performance degradation (and may involve costly pretraining) or a held-out subset of measurements for self-validation.
@@
In this work, we explore the possibility of modulating the frequency bias and hence the regularization effects of network priors in an architecture-agnostic manner, aiming to enhance a given architecture irrespective of its specific configurations. This is conceivable, in light of the recent body of theoretical and empirical evidence indicating that there are only a few key componenets within the architecture that are driving forces behind the spectral bias in DIP [30, 5, 41, 15]. Inspired by these findings, we develop simple yet effective methods from a frequency perspective to efficiently regularize the network priors, alleviating overfitting by curbing the over-fast convergence of high frequencies, all with minimal architectural modifications and computational costs. Specifically, we propose to (i) constrain the effective bandwidth of the white-noise input via simple spatial blurring or selected Fourier features, (ii) adjust the bandwidths of the interpolation-based upsamplers to control their attenuation (smoothing) extents, and (iii) regularize the Lipschitz constants of the layers to enforce function smoothness. We empirically find that our regularized network priors not only exhibit less vulnerability to overfitting but also tend to achieve better extrapolation and thus higher peak performance in inpainting-like tasks. Demonstrated in the context of MRI reconstruction, our methods significantly improve models across various architectural configurations, largely reducing the need for extensive architectural tuning (Fig. 1). The efficacy of our methods is also showcased in image inpainting and denoising tasks, where the control over frequency is adjusted accordingly. By minimizing architectural influences, our approach additionally offers a unique advantage in efficiency: a smaller, previously underperforming network, can now achieve on-par performance or even surpass a larger, heavily parameterized high-performing network. Our contributions are three-fold:
@@
  • We propose efficient methods that directly modulate the frequency bias in the untrained network priors in an architecture-agnostic manner, effectively addressing challenges related to architecture, overfitting, and runtime in a unified approach.
@@
  • The enhanced untrained networks match the performance of leading self-supervised methods and surpass the supervised method especially on out-of-domain data, with up to 86×86$\times$86 × reduction in runtime ($\sim 5$ mins/slice) for MRI reconstruction, serving as a stronger zero-shot reconstruction baseline than the original DIP scheme.
@@
  • Our empirical findings and results on medical and natural image reconstruction tasks shed light on the spectral behaviors of network-based priors—CNNs operating in a single-instance generative setting.
@@ 2 Related Work
Function Smoothness and Spectral Bias. Function smoothness, also referred to as function frequency, quantifies how much the output of a function varies with changes in its input [12]. Spectral bias [38, 47] is an implicit bias that favors learning functions changing at a slow rate (low-frequency), e.g., functions with a small Lipschitz constant. In visual domains, this is evident in the network’s output lacking subtle details. Many regularization techniques shown to aid generalization encourage smoothness implicitly, such as early stopping, $\ell_{2}$ regularizer [39]. Smoothness has been widely used as a model complexity measure in place of model size to account for the well-known "double-descent" phenomenon associated with generalization [34]. To explicitly promote smoothness, it is natural to penalize the norm of the input-output Jacobian [36, 17]. However, due to the high dimensionality of the output such as in accelerated MRI, computation of the Jacobian matrix during training is often intractable. Another efficient and prevalent solution is to constrain the network to be c-Lipschitz with a pre-defined Lipschitz constant c [33, 13]. We develop a suitable form of Lipschitz regularization for untrained networks by instead penalizing learned Lipschitz constants, with a novel aim of achieving architecture-agnostic untrained image reconstruction.
@@
Input Frequency and Generalization. Input has played an important role in helping neural networks represent signals of various frequencies. As in implicit neural representations (INRs) [42] and neural radiance fields (NeRFs) [32] where coordinates are mapped to RGB values, naively training with raw coordinates as inputs results in over-smoothing; encoding the input coordinates with sinusoidal functions of higher frequencies enables the network to represent higher frequencies [32, 43]. Rahaman et al. [38] also show theoretically and empirically that fitting becomes easier for the network when the input itself contains high-frequency components. However, it has recently been reported that the high-frequency positional input encodings lead to the failure of NeRFs in few-shot settings due to over-fitting [49]. Here, we show that similar issue also applies to untrained network priors and can be efficiently addressed by our methods.
Avoid Overfitting in Untrained Networks. A line of effort has been exclusively devoted to preventing overfitting to noisy images or measurements. Wang et al. [46] propose to track the running variance of the output for an early-stopping criterion, but it is found to be unstable in medical image reconstruction [2]. Yaman et al. [48] propose to split the available measurements into a training and a validation subset and use the latter for self-validation for automated early stopping. Transfer-learning-based untrained networks perform pre-training on synthetic data followed by fine-tuning [3, 35] or subspace optimization [2]. These methods aim to use fewer trainable parameters to avoid overfitting with little performance degradation. In contrast, our methods alleviate overfitting from a frequency perspective and enable significantly better performance while maintaining the same or less model-wise and computation-wise complexity.
Figure 2: Overview of the proposed regularized network priors for MRI reconstruction.
@@ 3.1 Preliminaries
Accelerated MRI The goal of accelerated MRI reconstruction is to recover a desired image $\mathbf{x\in\mathbb{C}}^{n}$ ($n=n_{h}\times n_{w}$) from a set of under-sampled k-space measurements. We focus on a multi-coil scheme in which the forward model is defined as:
$\mathbf{y}_{i}=\mathbf{A}_{i}\mathbf{x}+\epsilon\text{ with }\mathbf{A}_{i}=\mathbf{MFS}_{i},\quad i=1,\dots,c,$ (1)
@@
Untrained MRI Reconstruction is often framed as an inpainting problem where the network recovers the unacquired $k$-space measurements (masked) based on the acquired $k$-space data (observed). The image $\mathbf{x}$ is parameterized via a neural network $\mathbf{G_{\theta}(z)}$ with a fixed noise input vector $\mathbf{z}$ drawn from a uniform distribution $z\sim\mathcal{U}(0,1)$. With the aforementioned MRI forward model, the untrained network solves the following optimization problem:
@@
Such parameterization allows novel image priors to be designed dependent on the network architecture and the associated parameters. Nevertheless, many studies augment the untrained networks with traditional image regularizers [29], e.g., Total variation (TV), though it can only partially alleviate over-fitting [35, 2]. In our experimental sections, we show that TV is not as effective as ours in improving under-performing architectures.
@@ 3.2 Architecture-agnostic Frequency Regularization
To modulate the regularization effects of the network prior, we identify three sources of spectral bias within the network and introduce corresponding regulation methods, as depicted in Fig. 2. Each method requires only minimal hyperparameters to achieve architecture-agnostic control over the network priors.
@@
Figure 3: (a) Visualization of the 1D white noise, the low-passed noise via Gaussian blur, and Fourier features of various frequencies in the frequency domain. (b) Gaussian blurred noise input and Fourier features with lower $f_{c}$ ($L=4$ or $8$) are both effective in alleviating overfitting and enhancing the peak performance.
@@
Bandwidth-constrained Input. An important aspect of untrained networks that can be easily overlooked is their inputs. Conventionally, the inputs are randomly sampled from either a uniform or Gaussian distribution and are then mapped to an image. From a frequency perspective, such white-noise input comprises all frequencies with uniform intensities [11]. With this view in mind, we draw an analogy between untrained networks and INRs that map Fourier features to RGB values. Fourier features are sinusoid functions of the input coordinates $\mathrm{\mathbf{p}}$, i.e., $[\sin(\mathrm{2^{0}\pi\mathbf{p}}),\cos(\mathrm{2^{0}\pi\mathbf{p}}),...,\sin(2^{L-1}\pi\mathbf{\mathrm{\mathbf{p}}}),\cos(2^{L-1}\pi\mathbf{\mathrm{\mathbf{p}}})]$, where a larger $L$ assists the network in representing higher-frequency functions [43].
@@
Figure 4: Narrowing input’s bandwidth enhances performance as a "free lunch". The output becomes smoother as $\sigma$ increases, up to a certain point.
White noise with variance $\sigma^{2}$ exhibits an autocorrelation that is a scaled Dirac $\delta$-function $\sigma^{2}\delta(t)$, whose Fourier transform $\mathcal{F}$ has a constant magnitude $\sigma^{2}$ spanning all frequencies $\mu$, i.e., $\mathcal{F}\{\sigma^{2}\delta(t)\}(\mu)=\sigma^{2}$, $\mu\in\mathbb{R}$.
@@
In this sense, an untrained network can be viewed as mapping a broad spectrum of Fourier features to a target image (Fig. 3 (a)). We hypothesize that this enhances the network’s representation ability but likely incurs over-fitting due to the faster convergence of high-frequency components. To validate this, we applied a Gaussian blur filter $\mathcal{G}_{s,\sigma}$ on the white-noise input $z$ to remove a certain amount of high frequencies before passing it to the network, defined as $z*\mathcal{G}_{s,\sigma}$, where *** denotes convolution. The filter size $s$ and sigma value $\sigma$ that controls the filter’s bandwidth are hyperparameters. As exemplified in Fig. 4, simply adjusting $\sigma$ already brings significant gains without architectural changes. Similarly, substituting the noise input with Fourier features, with a carefully selected maximum frequency $f_{c}\propto L$ to limit the input’s effective bandwidth has also proven effective in Fig. 3 (b) and our ablation study (Tab. 23), further validating our hypothesis. Fourier features introduce frequency-diverse input akin to white noise but in a controlled manner, enabling frequency regularization.
Drawing from these empirical findings, we propose to regularize the input’s frequency range, achievable via simple spatial blurring or selected Fourier features, to efficiently mitigate overfitting and enhance network performance. While Gaussian blur and Fourier Features both offer bandwidth control, they yield different frequency representations (e.g., Fourier features are more discrete than Gaussian blurred noise), which may affect performance and susceptibility to parameter tuning differently. We present and compare both results in the experiments.
We note that inputs have been deemed crucial in guiding the generation of the final image not only in INRs but also in generative networks, e.g., transformed Fourier features have been used as an alternative to random input in StyleGAN3 [21], for a different goal of improving the architecture’s equivariance. In contrast, we use a low-passed version of Fourier features or noise input to prevent over-fast convergence on high frequencies and enable better extrapolation. Similar in spirit, recent work by [49] also shows that masking the high-frequency Fourier features helps NeRFs generalize in few-shot settings.
@@
Smoothness-controlled Upsampling. We found that constraining the input’s bandwidth significantly enhances shallower architectures, yet the improvement diminishes as the network depth increases. This could partly be attributable to the fact that subsequent layers along with nonlinearity can always generate new arbitrarily high frequencies, as theoretically and empirically proved in [38, 21].
We notice that the interpolation-based upsampling methods within the network, such as nearest neighbor and bilinear, are essentially implicit low-pass filters smoothing out the alias frequencies caused by the increased sampling rate. Prior works [30, 15] on image denoising have shown that these non-trainable upsampling methods are driving forces behind the spectral bias of the network priors, delaying the convergence of higher frequencies. Different upsamplers bias the network towards different spectral properties [30]. We confirm that upsampling also substantially influences image reconstruction and further reveal that an appropriate upsampling method can bridge the performance gap between two distinct architectures (Tab. 1).
Motivated by these results, we introduce a controllable upsampler to modulate the network’s spectral bias, especially the deeper architectures. We achieve this by first interleaving the input feature maps with zeros, and then convolving them with a customized low-pass filter with adjustable bandwidth (Fig. 2). For filter design, we adopt the Kaiser-Bessel window as it offers explicit control over the tradeoffs between passband ripple and stopband attenuation. The Kaiser window is defined as
@@
where $M$ is the desired spatial extent of the window, $\beta\geq 0$ is the shape parameter—the higher it is, the greater the stopband attenuation is (and generally the smoother the image is), and $I_{0}$ is the zeroth-order modified Bessel function of the first kind [20]. This plug-and-play upsampler can be inserted in different layers with different $M$ and $\beta$ hyperparameters, offering flexible and precise control.
Lipschitz-regularized Layers. Spectral bias towards low frequencies favors functions that do not change at a high rate, i.e., functions with small Lipschitz constants. A function $f:\mathcal{X}\rightarrow\mathcal{Y}$ is said to be Lipschitz continuous if there is a constant $k>0$ such that
$\|f(\mathrm{x_{1}})-f(\mathrm{x_{2}})\|_{\text{p}}\leq k\|\mathrm{x_{1}}-\mathrm{x_{2}}\|_{\text{p}}\quad\forall\mathrm{x_{1}},\mathrm{x_{2}}\in\mathcal{X},$ (4)
where $k$ is the Lipschitz constant that bounds how fast $f$ can change globally w.r.t. input perturbations.
@@
The Lipschitz constant of a convolutional layer is bounded by the operator norm of its weight matrix [13]. To bound a convolutional layer to a specific Lipschitz constant $k$, the layer with $m$ input channels, $c$ output channels and kernels of size $w\times h$ is first reshaped to a 2-D matrix $W^{\ell}\in\mathbb{R}^{n\times cwh}$, and then normalized as:

### v3

# Architecture-Agnostic Untrained Network Priors for Image Reconstruction with Frequency Regularization

@@ Abstract
Untrained networks inspired by deep image priors have shown promising capabilities in recovering high-quality images from noisy or partial measurements without requiring training sets. Their success is widely attributed to implicit regularization due to the spectral bias of suitable network architectures. However, the application of such network-based priors often entails superfluous architectural decisions, risks of overfitting, and lengthy optimization processes, all of which hinder their practicality. To address these challenges, we propose efficient architecture-agnostic techniques to directly modulate the spectral bias of network priors: 1) bandwidth-constrained input, 2) bandwidth-controllable upsamplers, and 3) Lipschitz-regularized convolutional layers. We show that, with just a few lines of code, we can reduce overfitting in underperforming architectures and close performance gaps with high-performing counterparts, minimizing the need for extensive architecture tuning. This makes it possible to employ a more compact model to achieve performance similar or superior to larger models while reducing runtime. Demonstrated on inpainting-like MRI reconstruction task, our results signify for the first time that architectural biases, overfitting, and runtime issues of untrained network priors can be simultaneously addressed without architectural modifications. Our code is publicly available  [^1] https://github.com/YilinLiu97/Untrained-Recon.git.
@@ 1 Introduction
Magnetic resonance imaging (MRI) is a mainstream imaging tool for medical diagnosis. Reconstructing MR images from raw measurements involves data transformation from a Fourier spectrum in $k$-space to image space [11, 28]. Since acquiring full $k$-space measurements is time-consuming, under-sampled $k$-space data are often collected to reduce scan times. This makes accelerated MRI reconstruction an ill-posed inverse problem that conventionally requires handcrafted priors [30, 25] to mitigate the resulting aliasing artifacts. While supervised learning methods based on convolutional neural networks (CNNs) have enhanced reconstruction quality with fewer measurements, their training relies on paired under-sampled and fully-sampled measurements, which are expensive to acquire and can affect robustness and generalization across different acquisition protocols or anatomical variations [21, 22].
@@
Figure 1: Improving underperforming architectures (SSIM ($\uparrow$)). Turning the left to the right simply by replacing the white-noise input with selected Fourier features or low-pass filtering the noise input, which can be implemented with a few lines of code.
@@
Instead of requiring large-scale datasets for capturing prior statistics, untrained networks [12] inspired by deep image prior (DIP) [43] use only the corrupted or partial measurements and a task-specific forward operator. Reconstruction is regularized solely by the inductive biases of the network architectures, enabling zero-shot self-supervised reconstruction for various imaging inverse problems [48, 26, 36]. Concretely, a CNN, which parameterizes the unknown desired image, is optimized such that the output image, transformed by the forward operator, matches the acquired measurements. Such parameterization exhibits surprisingly high resistance to noise and corruption, which acts as a form of implicit regularization. Recent studies have attributed this property to CNN’s inherent spectral bias—the tendency to fit the low-frequency signals before the high-frequency signals (e.g., noise) [39, 4]. The choice of network architecture is shown to be critically relevant to such bias [4, 29, 1].
@@
Despite the great promise, obtaining favorable results with untrained network priors is contingent upon two critical factors: an optimal architecture specific to a task and an early-stopping strategy to prevent overfitting to noisy or partial measurements. Furthermore, optimizing on a per-image basis makes the reconstruction process domain-agnostic but inherently slow [46]. While these issues are intertwined, with overfitting and runtime issues exacerbated by inappropriate and over-parameterized architectures, most existing efforts tackle these challenges separately. For architectural design, existing methods rely on either handcrafting or utilizing neural architecture search techniques [12, 6, 1, 5, 24, 14]. However, the lack of consensus on architectural priors often leads to laborious search. Another line of work is dedicated to preventing overfitting through oracle early-stopping criterions [46, 44], subspace optimization [2] or pretraining then fine-tuning [3, 34]. These methods mostly use fewer trainable parameters or hold out a subset of measurements for self-validation, in the spirit of traditional strategies, and often trade-off accuracy or involve costly pre-training.
@@
In this work, we explore the possibility of modulating the frequency bias and hence the regularization effects of network priors in an architecture-agnostic manner, aiming to enhance the performance of a given architecture irrespective of its configuration specifics. This is conceivable in light of the recent body of theoretical and empirical evidence indicating that there are only a few key components (e.g., unlearnt upsampling) within the architecture that are the driving forces behind the spectral bias in DIP [29, 4, 39, 13]. Motivated by these findings, we develop efficient methods from a frequency perspective to effectively regularize the network priors, alleviating overfitting by curbing the overly rapid convergence of undesired high-frequency components, all with minimal architectural modifications and computational costs. Specifically, we propose to (1) constrain the effective bandwidth of the input via blurring or using Fourier features with a narrower frequency range, (2) adjust the bandwidths of the interpolation-based upsamplers with controllable attenuation (smoothing) extents, and (3) regularize the Lipschitz constants of the convolutional layers to enforce function smoothness.
We found empirically that by mitigating convergence to high-frequency components, our regularized network priors not only exhibit less vulnerability to overfitting but also tend to achieve better extrapolation capabilities in inpainting tasks. In the context of MRI reconstruction, which is essentially an inpainting task occurring in $k$-space, our methods significantly improve models across various architectural configurations without necessitating extensive architectural tuning (Fig. 1). Their efficacy is also showcased in denoising and inpainting for natural images. By minimizing architectural influences, our approach additionally offers a unique advantage in efficiency: a smaller, previously underperforming network, can now achieve performance on-par with or even surpasses a larger, heavily parameterized high-performing network. Our contribution is three-fold:
@@
  • We propose efficient techniques to directly modulate the frequency bias in untrained network priors, addressing architectural design, overfitting, and runtime challenges in a unified, architecture-agnostic manner.
@@
  • The enhanced untrained networks match leading self-supervised methods with up to 90$\times$ faster runtime (1 hr/slice to $\sim$5  min/slice) for MRI reconstruction and surpass supervised methods on out-of-domain data.
@@
  • Our findings on medical and natural image reconstruction reveal the spectral behaviors of CNNs in a single-instance generative setting.
@@ 2 Related Work
Spectral Bias and function smoothness. Function smoothness, also referred to as function frequency, quantifies how much the output of a function varies with changes in its input [9]. Spectral bias [37, 45] is an implicit bias that favors learning functions changing at a slow rate (low-frequency), e.g., functions with a small Lipschitz constant. In visual domains, this is evident from the lack of subtle details in network outputs. Many regularization techniques to aid generalization, such as early stopping and $\ell_{2}$ regularization [38, 33], implicitly encourage smoothness. To explicitly promote smoothness, it is natural to penalize the norm of the input-output Jacobian [35, 15]. However, computation of the Jacobian matrix for the high-dimensional MRI data during training is very expensive. Another efficient and prevalent solution is to constrain the network to be c-Lipschitz with a pre-defined Lipschitz constant c [32, 10]. We followed this line with a novel aim of achieving architecture-agnostic untrained image reconstruction.
@@
Input frequency and overfitting. The network input plays an important role in helping the neural network represent signals of various frequencies. Pioneering work on spectral bias [37] showed theoretically and empirically that fitting high frequencies becomes easier, provided that the data manifold itself contains high-frequency components (Sec. 4 in [37]). This has directly motivated implicit neural representations (INRs) [40] and neural radiance fields (NeRFs) [31] where coordinates are mapped to RGB values: naively training with raw coordinates as inputs results in over-smoothing; encoding the input coordinates with sinusoidal functions of higher frequencies enables the network to represent higher frequencies [31, 41]. However, it has recently been reported that the high-frequency bands of NeRF’s input encodings incur overfitting and lead to failure in few-shot settings [47]. We found similar issues in untrained networks and propose to constrain the input’s frequency range to counteract the convergence of high frequencies.
Architecture-induced spectral bias. Recent studies on the working mechanisms of DIP reveal that unlearnt upsampling, with the low-pass characteristics of its interpolation filter, is responsible for the regularizing effects of DIP [29, 13, 4]. Liu et al. [29] showed that the fixed upsampling operations readily bias the architecture towards producing low-frequency outputs, critically influencing both the peak PSNR and the starting point at which performance decays. The convolutional layer is another core element that exhibits stronger frequency selectivity compared to fully-connected layers and 1D layers [4, 39]. These findings motivated us to operate directly on these elements to achieve architecture-agnostic control.
Avoiding overfitting is a primary goal of unsupervised reconstruction where only noisy or partial measurements are available. Wang et al. [44] proposed an early-stopping (ES) criterion by tracking the running variance of the output, but it is found to be unstable in medical settings [2]. Yaman et al. [46] proposed to split the available measurements into training and validation subsets and use the latter for self-validation for automated early stopping. While ES prevents overfitting, it cannot enhance the capability of a given architecture like our approach. Transfer-learning based methods aim to use fewer non-trainable parameters by performing pre-training followed by fine-tuning [3, 34] or subspace optimization [2]. In contrast, our method directly modulates the spectral bias to mitigate the convergence of undesired high frequencies, which is found to also improve the networks’ extrapolation capabilities, all while maintaining the model complexity.
Figure 2: Overview of the proposed regularized network priors for MRI reconstruction. By adjusting the bandwidth of the input, the interpolation-based upsampling, and regularizing the convolutional layers, our approach enables more direct control over the spectral bias of architectures with various depths and widths.
@@ 3.1 Preliminaries
Accelerated MRI The goal of accelerated MRI reconstruction is to recover a desired image $\mathbf{x\in\mathbb{C}}^{n}$ ($n=n_{\text{h}}\times n_{\text{w}}$) from a set of under-sampled k-space measurements. We focus on a multi-coil scheme in which the forward model is defined as
$\mathbf{y}_{i}=\mathbf{A}_{i}\mathbf{x}+\epsilon,\quad\mathbf{A}_{i}=\mathbf{MFS}_{i},\quad i=1,\dots,c,$ (1)
@@
Untrained MRI Reconstruction is often framed as an inpainting problem where the network recovers the unacquired $k$-space measurements (masked) based on the acquired $k$-space data (observed). The image $\mathbf{x}$ is parameterized via a neural network $\mathbf{G_{\theta}(z)}$ with a fixed noise input vector $\mathbf{z}$ drawn from a uniform distribution $z\sim\mathcal{U}(0,1)$. With the MRI forward model in Eq. 1, the untrained network solves the following optimization problem:
@@
This parameterization enables the design of novel image priors based on the network architecture and its parameters, rather than on handcrafted image space priors. Nevertheless, many studies augment the untrained networks with traditional image regularizers [27], e.g., total variation (TV), which promotes piecewise constant images and can only partially alleviate over-fitting [34, 2].
@@ 3.2 Architecture-Agnostic Frequency Regularization
To modulate the regularization effects of the network prior, we identify three core elements within the framework of DIP that lead to spectral bias and introduce corresponding regulation methods, as depicted in Fig. 2.
@@
Figure 3: (a) Visualization of the 1D white noise, the low-passed noise via Gaussian blur, and Fourier features of various frequencies in the frequency domain. (b) Limiting the input’s bandwidth via either Gaussian blur or Fourier features with lower $f_{c}$ ($L=4$ or $8$) are both effective in alleviating overfitting and enhancing the peak performance.
@@
Bandwidth-Constrained Input. Inspired by implicit neural representations (INRs), we rethink the role of inputs in untrained networks for learning different frequencies. Conventionally, the inputs are randomly sampled from either a uniform or Gaussian distribution and are then mapped to image intensities. From a frequency perspective, such input comprises all frequencies with uniform intensities [8], as white noise with variance $\sigma^{2}$ exhibits an autocorrelation that is a scaled Dirac $\delta$-function $\sigma^{2}\delta(t)$, whose Fourier transform $\mathcal{F}$ has a constant magnitude $\sigma^{2}$ spanning all frequencies $\mu$, i.e., $\mathcal{F}\{\sigma^{2}\delta(t)\}(\mu)=\sigma^{2}$, $\mu\in\mathbb{R}$. With this view in mind, we draw an analogy between untrained networks and INRs that map Fourier features to RGB values. Fourier features are sinusoid functions of the input coordinates $\mathrm{\mathbf{p}}$, i.e., $[\sin(\mathrm{2^{0}\pi\mathbf{p}}),\cos(\mathrm{2^{0}\pi\mathbf{p}}),...,\sin(2^{L-1}\pi\mathbf{\mathrm{\mathbf{p}}}),\cos(2^{L-1}\pi\mathbf{\mathrm{\mathbf{p}}})]$, where a larger $L$ assists the network in representing higher frequencies [41]. This is consistent with [37] which indicates that the frequency magnitudes that can be expressed by the network increase with increasing frequency in the data manifold.
In this sense, an untrained network can be viewed as mapping a broad spectrum of Fourier features to a target image (Fig. 3 (a)). We hypothesize that this promotes rapid convergence of high frequencies, which is likely to bias the network towards high-frequency artifacts and hinder its ability to exploit
@@
Figure 4: Narrowing input’s bandwidth promotes low frequencies and enhances extrapolation capabilities as a "free lunch". The output becomes smoother as $\sigma$ increases, up to a certain point.
spatial information effectively for extrapolation.
@@
To validate this, we applied a Gaussian blur filter $\mathcal{G}_{s,\sigma}$ on the noise input $z$ to remove a certain amount of high frequencies before passing it to the network: $z*\mathcal{G}_{s,\sigma}$, where * denotes convolution. The filter size $s$ and sigma value $\sigma$ that controls the filter’s bandwidth are hyperparameters. As exemplified in Fig. 4, simply adjusting $\sigma$ already brings significant gains without architectural changes. Alternatively, as shown in Fig. 3 (b), substituting the noise input with Fourier features, with a carefully selected maximum frequency $f_{c}\propto L$ (e.g., $L=4$ or 8) to narrow the input’s effective bandwidth, is also effective. As $L$ increases, e.g., $L=16$, the frequency range of Fourier features input approximates that of the original noise input, and the performance deteriorates similarly, further supporting our hypothesis. Fourier features introduce frequency-diverse input akin to white noise but in a controlled manner, enabling regularization over the frequency selectivity of the network.
In light of this empirical evidence, we propose to limit the input’s bandwidth to mitigate the fitting of high-frequency components in untrained networks, achievable efficiently via blurring or Fourier features with a narrowed frequency range. Gaussian blur and Fourier Features offer flexible bandwidth control over the input through hyperparameters $\{s,\sigma\}$ and $L$, respectively, which allows for scaling under higher noise level/undersampling rates (examples of 8×8$\times$8 × undersampling in Fig. 8). We experimented with more sophisticated schedulers for the hyperparameters, but found that simply fixing them throughout the training yields superior performance.
@@
Bandwidth-Controllable Upsampling. We observed that only constraining the input’s bandwidth significantly enhances shallower architectures, yet the improvement diminishes as the network depth increases (Tab. 23). This could be partly attributed to the increased network layers that can always generate new arbitrarily high frequencies [37, 19].
We note that the interpolation-based upsampling methods within the network, such as nearest neighbor and bilinear, essentially act as implicit low-pass filters, attenuating the alias frequencies caused by the increased sampling rate of the input feature maps. Prior works [29, 13] in denoising have shown that these non-trainable upsampling methods are driving forces behind the spectral bias of DIP, delaying the convergence of higher frequencies. Different upsamplers bias the network towards different spectral properties, depending on the bandwidth of the interpolation filter [29]. Here, we show in Tab. 1 that upsampling also substantially influences the performance image reconstruction.
Motivated by these results, we introduce an upsampler with controllable bandwidth so as to modulate the network’s spectral bias, especially for deeper architectures. We construct it by 1) first interleaving the input feature maps with zeros, and then 2) convolving them with a customized low-pass filter with adjustable bandwidth (Fig. 2). For filter design, we adopt the Kaiser-Bessel window [18] as it offers explicit control over the tradeoffs between passband ripple and stopband attenuation. The Kaiser window is defined as
@@
where $M$ is the desired spatial extent of the window, $\beta\geq 0$ is the shape parameter—the higher it is, the greater the stopband attenuation is (and generally the smoother the image is), and $I_{0}$ is the zeroth-order modified Bessel function of the first kind. This plug-and-play upsampler can be inserted in different layers with different $M$ and $\beta$ hyperparameters, offering flexible and precise control.
Lipschitz-Regularized Layers. Compared to the non-trainable upsampling that only attenuates signals, the network layer with nonlinearies is the only operation capable of generating new frequencies [19]. We regularize their Lipschitz constants to control their sensitivity to input variations, which in turn can affect the spectral bias. Formally, a function $f:\mathcal{X}\rightarrow\mathcal{Y}$ is said to be Lipschitz continuous if there is a constant $k>0$ such that $\|f(\mathrm{x_{1}})-f(\mathrm{x_{2}})\|_{\text{p}}\leq k\|\mathrm{x_{1}}-\mathrm{x_{2}}\|_{\text{p}}\quad\forall\mathrm{x_{1}},\mathrm{x_{2}}\in\mathcal{X},$ where $k$ is the Lipschitz constant that bounds how fast $f$ can change globally w.r.t. input perturbations. Spectral bias towards low frequencies favors functions with small Lipchitz constants.
@@
The Lipschitz constant of a convolutional layer is bounded by the operator norm of its weight matrix [10]. To bound a convolutional layer to a specific Lipschitz constant $k$, the layer with $m$ input channels, $c$ output channels and kernels of size $w\times h$ is first reshaped to a 2-D matrix $W^{\ell}\in\mathbb{R}^{n\times cwh}$, and normalized as